INAF queue observing, 2018 September 24-25

Observer: R. Carini, A. Gargiulo
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg
Instrument: LUCI

Summary

This night is shared with Eng team. We take second half of the night, starting at ~00:00 local At the sunset the sky is cloudy
instruments: MODS2 red grating is not working
LBC-blue has some issues, LUCI1 works only if the guide star is on the right of the patrol field, LUCI1 AO doesn't work.

At 00:15 the sky is mostly clear but the humidity is at 100% and the temperature at the dew point. The dome is closed. At 00:40 the sky is covered

Night duration (12\xB0 el twilight) 2:30- 11:45 UT = 9 h 15m

Observing time: 0 0m (0%)

Weather time loss: 4h 45m (100%)

technical time loss m (%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Calibration Dark Cresci,Nisini,Nardini,Perna          
arcs and flats Cresci,Nisini field2          
flats ima Cresci, Nisini          

Details

calibration

Cresci dark spectra luci1.20180925.0003-0007, luci2.20180925.0007-0011 dit 150 ndit 1 mer norm (Ok also for Perna)
  • dark tell luci1.20180925.0008-0012, luci2.20180925.0012-0016 dit 3 ndit 3 lir int
  • dark ima QSO J0129 luci1.0013-0017, luci2.0017-0021 dit 2.51 ndit 20 lir int
  • dark ima QSO J2054, J2310 luci1.0018-0022, luci2.0022-0026 dit 2.51 ndit 5 lir int

Nisini dark ima luci1.0023-0027, luci2.0027-0031 dit 5 ndit 3 lir int
  • dark spectra luci1.0028-0032, luci2.0032-0036 dit 45 ndit 5 lir int
  • dark tell luci1.0033-0037, luci2.0037-0041 dit 2.5 ndit 25 * dark ima luci1.0038-0042, luci2.0042-0046 dit 2.7 ndit 3

Nardini dark tell luci1.0043-0047 dit 5 ndit 3 lir int
  • dark spectra luci1.0048-0052 dit 110 ndit 1 mer norm
  • dark spectra luci1.0053-0057 dit 100 ndit 1 mer norm

Perna dark tell luci1.0058-0062, luci2.0047-0051 dit 3 ndit 5 lir int (ok also for tell Cassar\xE1)

Cassar\xE1 dark spectra luci1.0063-0067, luci2.0052-0056 dit 200 ndit 1 mer norm

* Flats and arcs *

*Cresci luci1.0068-0072, luci2.0057-0061 dark
  • flats spectra ==luci1.0073-0077, luci2.0062-0066==zjspec
  • arc luci1.0078-0079, luci2.0067-0068 zjspec, Neon
  • arc luci1.0080-0081, luci2.0069-0070 zjspec, Argon
  • ==luci1.0082-0086, luci2.0071-0075==dark
  • flat ima luci1.0087-0091, luci2.0076-0080 J

Nisini field2
  • dark luci1.0092-0096, luci2.0081-0085
  • flats luci1.0097-0101, luci2.0086-0090 H
  • arc luci1.00102-0103, luci2.0091-0092 H, Ar
  • arc luci1.0104-0105, luci2.0093-0094 H, Xe * dark luci1.0106-0110, luci2.0095-0099
  • flats luci1.0111-0115, luci2.0100-0104 J
  • arc luci1.00116-0117, luci2.0105-0106 J, Ar
  • arc luci1.0118-0119, luci2.0107-0108 J, Xe
  • dark luci1.0120-0124, luci2.0109-0113
  • during the integration the GUI frozen
  • flat luci1.00125-0129, luci2.0114-0118 K
  • arc luci1.0130-0131, luci2.0119-0120 K, Ne
  • arc luci1.0132-0133, luci2.0121-0122 K, Ar
  • arc luci1.0134-0135, luci2.0123-0124 K, Ar

flat ima
  • dark luci1.0136-0140, luci2.0125-0129
  • FeII luci1.0141-0145, luci2.0130-0134
  • luci 2 doesn't change filter during the dark, we have to load again the script
  • error again: error commit setup
  • SA fix the problem, we finish the flats for imaging

-- %USERSIG{RobertaCarini - 2018-09-25}%

Comments

 
Topic revision: r5 - 01 Oct 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback